OpenStudy (anonymous):

the inverse of 3x+4/2x-3

OpenStudy (anonymous):

solve \[x=\frac{3y+4}{2y-3}\] for y you need the steps?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did it, but i'm not sure of it

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i think this function is its own inverse

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i can write out the steps if you like

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[x=\frac{3y+4}{2y-3}\] \[x(2y-3)=3y+4\] \[2xy-3x=3y+4\] \[2xy-3y=3x+4\] \[y(2x-3)=3x+4\] \[y=\frac{3x+4}{2x-3}\] same thing

OpenStudy (anonymous):

yeah :d thks alot